• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 07:57:55 on July26,2005

    Entry number 149503

    keyword=Owl Shift Summary

    Owl Shift Summary
    ----------------------
    Run 3826 - production run at 40 uA
    Run 3827 - production run at 40 uA *20 minutes
    
    Beam taken away by MCC to try and fix pathlength - to deliver better beam to 
    Hall C, and hopfully better quality beam for the Compton in Hall A.
    
    Estimated time for completion is 2:15.
    
    2:30 - MCC needs more time for pathlength adjustment.  They hope to finish 
    around 3:00 or 3:30.  See halog 149481
    .
    
    3:25 - PSS ioc went down.  MCC has to fix this before they can continue 
    pathlength work.  Probably down for another hour or so.
    
    4:00 - MCC is having some software problem that won't allow them to clear 
    faults in order to deliver beam.  They will let us know when it gets fixed.  No 
    time estimate.
    
    6:00 MCC software problem fixed.
    
    7:18 - update from MCC.  They are working on adjusting the pathlength now.
    
    Shift ended with no beam.
    
    Run 3828 - pedestal run
    Run 3829 - pedestal run
    Run 3830 - pedestal run
    Run 3831 - pedestal run
    Run 3832 - pedestal run
    Run 3833 - pedestal run
    Run 3834 - pedestal run
    Run 3835 - pedestal run